40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Mon Sep 29 23:28:37 2014, Evan, DailyProgress, PMC, Re-inserted PMC, moved BB EOM noiseBudget.pdf
    Reply  Tue Sep 30 21:45:14 2014, Evan, DailyProgress, PMC, Re-inserted PMC, moved BB EOM 
    Reply  Mon Dec 15 20:43:49 2014, Aidan, Kate, Evan, DailyProgress, PMC, South Faraday isolator needs help 
       Reply  Sat Jun 20 22:59:58 2015, Evan, DailyProgress, optic, South Faraday isolator path reworked southfi.jpg
Message ID: 1534     Entry time: Mon Dec 15 20:43:49 2014     In reply to: 1528     Reply to this: 1551
Author: Aidan, Kate, Evan 
Type: DailyProgress 
Category: PMC 
Subject: South Faraday isolator needs help 

Quote:

The laser beam entering the first Faraday isolator appears to be 1–2 mm too low. It is clipping on the input aperture, and the transmitted beam looks like crap.

When Aidan and I turned on the south laser today, we found that the transmitted beam out of this Faraday was entirely crap. It was blindingly obvious on an IR card, and only 50 uW was making it to the input of the PMC. The rest was scattering at wide angles at the Faraday output port.

It is not clear to me how the pointing through the Faraday could have deteriorated, since it is on a solid metal mount and is only 10 cm from the output of the laser.

At any rate, I was able to "recover" the previous performance (i.e., crappy but workable) by placing the Faraday isolator slightly further down in the optical path. Before, the layout was:

Laser -> QWP -> HWP -> Faraday -> lens -> HWP -> steering mirror -> PMC EOM,

and the HWP angles were -1 deg and 167 deg, respectively.  Now the layout is

Laser -> QWP -> HWP -> lens -> steering mirror -> Faraday -> HWP -> PMC EOM,

and the HWP angles are 341 deg and 167 deg. The first HWP angle is chosen so that 20 mW is transmitted through the Faraday (the rest is dumped at the Faraday's various output ports). The second HWP angle is chosen to send s polarization through the PMC EOM. I then had to resteer through the PMC EOM and through the PMC. With 20 mW incident on the PMC, the transmission is 11 mW. Not great, but about the same as the previous situation.

I remark that the south optical path between the laser and the PMC should be reworked as soon as is feasible, because what I've done is a hack job to keep things moving. Either the Faraday mount needs to be remachined, or the optical path needs to be redesigned to allow for proper steering through the Faraday. Additionally, the table surface next to the laser mounts is noticeably warm to the touch, so I do not recommend trying to shim up the laser (as it may negatively impact the heatsinking).

ELOG V3.1.3-